How many grams are in 5.60 moles of SO2?


1
Expert's answer
2021-05-07T05:25:06-0400

Molar mass of "SO_2=64g"


No. of moles "=\dfrac{\text{ Mass of } SO_2}{\text{Molar mass of } SO_2}"


"\Rightarrow 5.60=\dfrac{x}{64}\Rightarrow x=64\times 5.6=358.4g"


Hence The mass of "SO_2" required is 358.4g